# Green functions, Hitchin’s formula and curvature equations on tori

### v1

@@ 3.2. Examples that evey number in $\{4,6,8,10\}$ occurs
We turn to the proof of Theorem 1.4 (1)-(3) (the proof of Theorem 1.4 (3-1)-(3-2) will be postponed in Section 5). Recall the following result from [13].
@@
(3.15) $\det D^{2}G\Big{(}\pm q;\frac{\tau}{2}\Big{)}>0,\quad\text{and}\quad\det D^{2}G\Big{(}z;\frac{\tau}{2}\Big{)}<0\;\text{for}\;z\in\Big{\{}\frac{1}{2},\frac{\tau}{4},\frac{1}{2}+\frac{\tau}{4}\Big{\}}.$
@@
Similarly, we can prove the following result as a counterpart of Theorem 3.5 for those $\tau$ such that $G(z)=G(z;\tau)$ has $5$ critical points.
Theorem 3.7.
Proof.
Fix any $\tau$ such that $G(z)=G(z;\tau)$ has $5$ critical points $\frac{\omega_{k}}{2}$, $k=1,2,3$, and $\pm q\notin E_{\tau}[2]$. Again these critical points are all non-degenerate and
Assume by contradiction that there is $p_{n}\to 0$ such that $G_{p_{n}}(z)$ has more than $6$ critical points, i.e. has another critical point $z_{p_{n}}\notin E_{\tau}[2]\cup\{\pm q_{p_{n}}\}$. Then $z_{p_{n}}$ is a nontrivial critical point. Up to a subsequence, we may assume $z_{p_{n}}\to z_{0}\in E_{\tau}$. Then it was proved in [12, Theorem 1.5] that $z_{0}\neq 0$ in $E_{\tau}$, so $z_{0}$ is a critical point of $G(z)$, i.e. $z_{0}=\frac{\omega_{k}}{2}$ for some $k=1,2,3$ or $z_{0}=\pm q$. But this is a contradiction with the fact that the critical point of $G_{p_{n}}(z)$ in the small neighborhood of $z_{0}$ is unique. Therefore, by taking $\varepsilon>0$ smaller if necessary, we conclude that for any $0<|p|<\varepsilon$, $G_{p}(z)$ exactly $6$ critical points that are all non-degenerate, with $\det D^{2}G_{p}(\frac{\omega_{k}}{2})<0$ for $k\in\{0,1,2,3\}$ and $\det D^{2}G_{p}(\pm q_{p})>0$, namely $\pm q_{p}$ are minimal points of $G_{p}(z)$. This is the case in Corollary 2.10-(1) with $6$ critical points.
@@ 4. Generic non-degeneracy of critical points
Remark 4.1.
In [17], Hitchin’s formula (1.8) was used as a function of $\tau$ with fixed $(r,s)$. Here we use Hitchin’s formula (1.8) in a different way, i.e. we fix $\tau$ and consider it as a function of $(r,s)$.
@@
To prove Theorem 1.8, we need to prove the following result.
@@
is of Lebesgue measure zero, so $\mathcal{O}_{\tau}$ is also of Lebesgue measure zero.
@@
Denote by $q_{1},\cdots,q_{N}$ the $N$ critical points of $G_{p_{0}}(z)$. Since they are all non-degenerate, it follows from the implicit function theorem that there is $\varepsilon>0$ small such that for any $0<|p-p_{0}|<\varepsilon$, $G_{p}(z)$ has $N$ critical points $q_{p,1},\cdots,q_{p,N}$ (in small neighborhoods of each critical point $q_{j}$ of $G_{p_{0}}(z)$) and they are also non-degenerate. Assume by contradiction that there is $p_{n}\to p_{0}$ such that $G_{p_{n}}(z)$ has more than $N$ critcal points, i.e. has another nontrivial critical point $q_{p_{n},N+1}\in E_{\tau}\setminus\{q_{p_{n},j}\}_{j=1}^{N}$. Up to a subsequence, we may assume $q_{p_{n},N+1}\to q$. Write $q_{p_{n},N+1}=r_{n}+s_{n}\tau$ with $\mathbb{R}^{2}\setminus\frac{1}{2}\mathbb{Z}^{2}\ni(r_{n},s_{n})\to(r,s)$ and $q=r+s\tau$. Then (2.1) implies
@@ 5.2. Consequences
$\Xi_{1,\infty}:=\mathbb{C}\setminus(\cup_{k=0}^{3}\overline{\mathcal{B}_{k}}\cup\{e_{k}:k=1,2,3\}),$
$\Xi_{2,\infty}:=\mathcal{B}_{3}\setminus(\cup_{k=0}^{2}\overline{\mathcal{B}_{k}}\cup\{e_{k}:k=1,2,3\}).$
Then $\Xi_{j,\infty}$ are both unbounded, $G_{p}(z)$ has at least $6$ critical points for any $\wp(p)\in\Xi_{1,\infty}$, and $G_{p}(z)$ has exactly either $4$ or $8$ critical points for almost all $\wp(p)\in\Xi_{2,\infty}$.
@@ 5.3. Two special examples
    $\Xi_{4}:=\mathcal{B}_{2}\setminus\overline{\mathcal{B}_{3}}\neq\emptyset,\quad\widetilde{\Xi}_{4}:=\mathcal{B}_{1}\setminus\overline{\mathcal{B}_{3}}=-\Xi_{2},$
@@ 6. Connection with the curvature equation
(6.10) $y^{\prime\prime}(z)=\Big{[}\frac{3}{4}(\wp(z+p)+\wp(z-p))+A(\zeta(z+p)-\zeta(z-p))+B\Big{]}y(z),\;z\in\mathbb{C}$
@@
Now $q_{p}\to q\notin E_{\tau}[2]$ as $p\to 0$ implies that $f_{p}(z)\to f_{0}(z)=\exp(2\zeta(q)z)\frac{\sigma(z-q)}{\sigma(z+q)}$ (Note from (6.2) that $f_{0}(z)\equiv-1$ if $q=\frac{\omega_{k}}{2}$ for some $k=1,2,3$) and $u_{p}(z)\to u(z)=\log\frac{8|f_{0}^{\prime}(z)|^{2}}{(1+|f_{0}(z)|^{2})^{2}}$, where $u(z)$ is an even solution of (1.2).

### v2

@@ 3.2. Examples that evey number in $\{4,6,8,10\}$ occurs
Here we can prove the following result as a counterpart of Theorem 3.5 for those $\tau$ such that $G(z)=G(z;\tau)$ has $5$ critical points.
Theorem 3.6.
Proof.
Fix any $\tau$ such that $G(z)=G(z;\tau)$ has $5$ critical points $\frac{\omega_{k}}{2}$, $k=1,2,3$, and $\pm q\notin E_{\tau}[2]$. It follows from Theorems B and 1.2 that these critical points are all non-degenerate and
Assume by contradiction that there is $p_{n}\to 0$ such that $G_{p_{n}}(z)$ has more than $6$ critical points, i.e. has another critical point $z_{p_{n}}\notin E_{\tau}[2]\cup\{\pm q_{p_{n}}\}$. Then $z_{p_{n}}$ is a nontrivial critical point. Up to a subsequence, we may assume $z_{p_{n}}\to z_{0}\in E_{\tau}$. Suppose $z_{0}=0$ in $E_{\tau}$, by replacing $z_{p_{n}}$ with some element from $z_{p_{n}}+\mathbb{Z}+\mathbb{Z}\tau$, we may assume $z_{p_{n}}=r_{n}+s_{n}\tau\to 0$ with $\mathbb{R}\ni r_{n},s_{n}\to 0$. Then it follows from the Laurent series of $\zeta(z)$ and $\wp(z)$ that
$\zeta(z_{p_{n}})=\frac{1}{z_{p_{n}}}+O(|z_{p_{n}}|^{3}),\quad\wp(z_{p_{n}})=\frac{1}{z_{p_{n}}^{2}}+O(|z_{p_{n}}|^{2}),$
$\wp^{\prime}(z_{p_{n}})=\frac{-2}{z_{p_{n}}^{3}}+O(|z_{p_{n}}|).$
Inserting these into (1.8), we obtain
$\wp(p_{n})=\wp(z_{p_{n}})+\frac{\wp^{\prime}(z_{p_{n}})}{2(\zeta(z_{p_{n}})-r_{n}\eta_{1}-s_{n}\eta_{2})}=-\frac{r_{n}\eta_{1}+s_{n}\eta_{2}}{z_{p_{n}}}+O(|z_{p_{n}}|^{2}).$
Note from $\operatorname{Im}\tau>0$ and $\mathbb{R}^{2}\ni(r_{n},s_{n})\to(0,0)$ that $\frac{r_{n}\eta_{1}+s_{n}\eta_{2}}{z_{p_{n}}}=\frac{r_{n}\eta_{1}+s_{n}\eta_{2}}{r_{n}+s_{n}\tau}$ is uniformly bounded, we obtain a contradiction with $\wp(p_{n})\to\infty$ as $p_{n}\to 0$. This proves $z_{0}\neq 0$ in $E_{\tau}$, so $z_{0}$ is a critical point of $G(z)$, i.e. $z_{0}=\frac{\omega_{k}}{2}$ for some $k=1,2,3$ or $z_{0}=\pm q$. But this is a contradiction with the fact that the critical point of $G_{p_{n}}(z)$ in the small neighborhood of $z_{0}$ is unique. Therefore, by taking $\varepsilon>0$ smaller if necessary, we conclude that for any $0<|p|<\varepsilon$, $G_{p}(z)$ exactly $6$ critical points that are all non-degenerate, with $\det D^{2}G_{p}(\frac{\omega_{k}}{2})<0$ for $k\in\{0,1,2,3\}$ and $\det D^{2}G_{p}(\pm q_{p})>0$, namely $\pm q_{p}$ are minimal points of $G_{p}(z)$. This is the case in Corollary 2.10-(1) with $6$ critical points.
Now we turn to the proof of Theorem 1.4 (1)-(3) (the proof of Theorem 1.4 (3-1)-(3-2) will be postponed in Section 5). Recall the following result from [13].
@@
(3.15) $\det D^{2}G\Big(\pm q;\frac{\tau}{2}\Big)>0,\;\text{and}\;\det D^{2}G\Big(z;\frac{\tau}{2}\Big)<0\;\text{for}\;z\in\Big\{\frac{1}{2},\frac{\tau}{4},\frac{1}{2}+\frac{\tau}{4}\Big\}.$
@@ 4. Generic non-degeneracy of critical points
Remark that in [17], Hitchin’s formula (1.8) was used as a function of $\tau$ with fixed $(r,s)$. Here we use Hitchin’s formula (1.8) in a different way, i.e. we fix $\tau$ and consider it as a function of $(r,s)$.
@@
To prove Theorem 1.8, we need to establish a precise formula of the Hessian $\det D^{2}G_{p}(q)$ in terms of $f$.
@@
is of Lebesgue measure zero, so $\mathcal{O}_{\tau}$ is also of Lebesgue measure zero. This proves that $\mathcal{O}_{\tau}\cup\mathcal{T}_{\tau}$ is of Lebesgue measure zero, and all critical points of $G_{p}(z)$ are non-degenerate for any $p\in E_{\tau}\setminus(E_{\tau}[2]\cup\mathcal{O}_{\tau}\cup\mathcal{T}_{\tau})$.
@@
Denote by $q_{1},\cdots,q_{N}$ the $N$ critical points of $G_{p_{0}}(z)$. Since they are all non-degenerate, it follows from the implicit function theorem that there is $\varepsilon>0$ small such that for any $0<|p-p_{0}|<\varepsilon$, $G_{p}(z)$ has $N$ critical points $q_{p,1},\cdots,q_{p,N}$ (in small neighborhoods of each critical point $q_{j}$ of $G_{p_{0}}(z)$) and they are also non-degenerate. Assume by contradiction that there is $p_{n}\to p_{0}$ such that $G_{p_{n}}(z)$ has more than $N$ critical points, i.e. has another nontrivial critical point $q_{p_{n},N+1}\in E_{\tau}\setminus\{q_{p_{n},j}\}_{j=1}^{N}$. Up to a subsequence, we may assume $q_{p_{n},N+1}\to q$. Write $q_{p_{n},N+1}=r_{n}+s_{n}\tau$ with $\mathbb{R}^{2}\setminus\frac{1}{2}\mathbb{Z}^{2}\ni(r_{n},s_{n})\to(r,s)$ and $q=r+s\tau$. Then (2.1) implies
@@ 5. Counting the number of critical points
$\Xi_{1,\infty}:=\mathbb{C}\setminus\big(\{e_{1},e_{2},e_{3}\}\cup\cup_{k=0}^{3}\overline{\mathcal{B}_{k}}\big),$
$\Xi_{2,\infty}:=\mathcal{B}_{3}\setminus\big(\{e_{1},e_{2},e_{3}\}\cup\cup_{k=0}^{2}\overline{\mathcal{B}_{k}}\big).$
Then $\Xi_{j,\infty}$ are both unbounded, $G_{p}(z)$ has at least $6$ critical points for any $\wp(p)\in\Xi_{1,\infty}$, $G_{p}(z)$ has exactly either $6$ or $10$ critical points for almost all $\wp(p)\in\Xi_{1,\infty}$, and $G_{p}(z)$ has exactly either $4$ or $8$ critical points for almost all $\wp(p)\in\Xi_{2,\infty}$.
@@ 5.3. Two special examples
    $\Xi_{4}:=\mathcal{B}_{2}\setminus\overline{\mathcal{B}_{3}}\neq\emptyset,\quad\widetilde{\Xi}_{4}:=\mathcal{B}_{1}\setminus\overline{\mathcal{B}_{3}}=-\Xi_{4},$
@@ 6. Connection with the curvature equation
(6.10) $y^{\prime\prime}(z)=\Big[$ $\frac{3}{4}(\wp(z+p)+\wp(z-p))$
$+A(\zeta(z+p)-\zeta(z-p))+B\Big]y(z),\quad z\in\mathbb{C}$
@@
Now $q_{p}\to q\notin E_{\tau}[2]$ as $p\to 0$ implies that $f_{p}(z)\to f_{0}(z)=\exp(2\zeta(q)z)\frac{\sigma(z-q)}{\sigma(z+q)}$ and $u_{p}(z)\to u(z)=\log\frac{8|f_{0}^{\prime}(z)|^{2}}{(1+|f_{0}(z)|^{2})^{2}}$, where $u(z)$ is an even solution of (1.2) (Note that $q\notin E_{\tau}[2]$ is crucial here, because it follows from (6.2) that $f_{0}(z)\equiv-1$ and $u_{p}(z)\to$-$\infty$ locally if $q=\frac{\omega_{k}}{2}$ for some $k=1,2,3$).